Seth Fitch owns a small retail ice cream parlor. He is considering expanding the business and has identified two attractive alternatives. One involves purchasing a machine that would enable Mr. Fitch to offer frozen yogurt to customers. The machine would cost $7,530 and has an expected useful life of three years with no salvage value. Additional annual cash revenues and cash operating expenses associated with selling yogurt are expected to be $5,950 and $880, respectively. Alternatively, Mr. Fitch could purchase for $9,720 the equipment necessary to serve cappuccino. That equipment has an expected useful life of four years and no salvage value. Additional annual cash revenues and cash operating expenses associated with selling cappuccino are expected to be $8,380 and $2,410, respectively. Income before taxes earned by the ice cream parlor is taxed at an effective rate of 20 percent. Required Determine the payback period and unadjusted rate of return (use average investment) for each alternative. (Round your answers to 2 decimal places.)
Answer #1

Alternative 1:

Cash flows after taxes = [(5950-880)*0.80+(7530/3)*0.20]

= $4558

Payback period = 7530/4558

= 1.65 years

Net income = (5950-880-2510)*80% Note : 7530/3 = 2510

= $2048

Unadjusted rate of return = 2048/3765 Note : 7530/2 = 3765

= 54.39%

Alternative 2:

Annual cash flows after taxes = [(8380-2410)*0.80+(9720/4)*0.2]

= $5262

Payback period = 9720/5350

= 1.81 years

Net income = (8380 - 2410 - 2430) *80%

= $2832

unadjusted rate of return = 2832/ 4860 Note : 9720/2 = 4860

= 58.27%